OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (mayankdevnani):

like 2x+2=6 2x=4 x=2

OpenStudy (mayankdevnani):

so x is variable and 2 is solution

OpenStudy (mayankdevnani):

ok @rhondaycole

OpenStudy (anonymous):

consider the equation 2x-3=0 now what should i put in x such that RHS become equall to LHS so x= 3/2 is the solution if you put x=3/2 in the above equation then the equation is satisfied.

OpenStudy (anonymous):

thus the value of a variable in an equation that satisfies that equation is called a solution to that equation